OpenStudy (anonymous):

slope 2; through (5-2)

OpenStudy (anonymous):

What do u want me to do?

OpenStudy (abmon98):

y+2=2(x-5), y=2x-12

OpenStudy (abmon98):

i used y-y1=m(x-x1)

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ 5}}\quad ,&{\color{blue}{ -2}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= 2 \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form} \)